OpenStudy (anonymous):

Grace is making party hats for her upcoming birthday party. A single hat has a diameter of 15 centimeters and a slant height of 28 centimeters. What is the lateral area of each party hat? 836.45 cm2 659.73 cm2 2,026.33 cm2 1,319.47 cm2

OpenStudy (anonymous):

I think Lateral Area of Cone (Hat is having a shape of Cone Only) is given by: \[Lateral \; Area = \frac{1}{2} \times \pi \times Diameter \times Slant \; Height\]

OpenStudy (anonymous):

Does that work???

OpenStudy (anonymous):

one sec

OpenStudy (anonymous):

tring to figure out what all to plug in and were

OpenStudy (anonymous):

Diameter = 15 Slant Height = 28 \(\pi = 3.14\) And then divide by 2 at last..

OpenStudy (anonymous):

i got 329.7

OpenStudy (anonymous):

Calculation mistake..

OpenStudy (anonymous):

Tell me the value you got for this: \[15 \times 28 \times 3.14 = ??\]

OpenStudy (anonymous):

1318.8

OpenStudy (anonymous):

659.4

OpenStudy (anonymous):

Yes now you are right..